Suppose you won $15 on a lotto ticket at the local 7-Eleven and decided to spend all the winnings on candy bars and bags of peanuts. Candy bars cost $0.75 each while bags of peanuts cost $1.50 each. a. Complete the table below (gray-shaded cells) showing the alternative combinations of the two products that are available. Instructions: Enter your answers as whole numbers. Alternative Combinations Goods 12 16 20 Candy bars Bags of peanuts ok b. Use the data in the table to plot a budget line in the graph. instructions: Use the tool provided Budget to draw a single budget line. Plot each of the six combinations (A - F) above.
Answer #1

No. of Candy Bars

Price ($)

Expenditure

Bags of Peanuts

Price ($)

Expenditure

Total Expenditure ($)

0

0.75

0*$0.75=0

$15/1.5 = 10

1.5

10*$1.5 = $15

15

4

0.75

4*$0.75=$3

($15-$3)/1.5 = 8

1.5

8*$1.5 = $12

15

8

0.75

8*$0.75=$6

($15-$6)/1.5 = 6

1.5

6*$1.5 = $9

15

12

0.75

12*$0.75=$9

($15-$9)/1.5 = 4

1.5

4*$1.5 = $6

15

16

0.75

16*$0.75=$12

($15-$12)/1.5 = 2

1.5

2*$1.5 = $3

15

20

0.75

20*$0.75=$15

($15-$15)/1.5 = 0

1.5

0

15

Final Answer

No. of Candy Bars

0

4

8

12

16

20

Bags of Peanuts

10

8

6

4

2

0

No of Peanuts 12 10 Series1 No. of candy bars 10 15 20 25 The horizontal axis is the candy bars and the vertical axis is peanuts. Budget line represents combination of two goods that can be purchased within the given level of budget No. of Candy Bars 0 Bags of Peanuts 10 12 8 6 16 2 20 0 8